Managerial Accounting

Can anyone Help me with these, and include the calculations in order for me not to hardcode the numbers because I need to use the excel formulas.

Thank you

Project #2 Sales Projections in Units January February March April |May 13,402 45,819 44,164 53,722 52,482 Projected Sales PrSALES BUDGET January February March April Projected Sales in Units Selling Price per Unit |Anticipated Sales DollarsSELLING & ADMINISTRATIVE EXPENSE BUDGET January February March Sales in Units Variable S&A Rate/Unit Anticipated Variable CosMATERIALS REQUIREMENTS BUDGET | January February March April Units to be Produced Material/Unit Total Production Need DesiredDIRECT LABOR BUDGET January February March Number of Units Produced Direct Labor Hours/Unit Total Direct Labor Hours Needed CMANUFACTURING OVERHEAD BUDGET January February March Production in Units Variable OH Rate/Unit Anticipated VOH Costs Fixed OHPRODUCT COST PER UNIT Item Amount Direct Materials Direct Labor Manufacturing Overhead Total Unit CostPROJECTED INCOME STATEMENT January February March Sales in Units Sales Revenue Cost of Goods Sold: Gross Profit (Margin) SellSENSITIVITY ANALYSIS (Print and submit AFTER submitting budget online) Situation #1: What would the effect be on the ProjecteProject #2 - Budgeting Learning Objectives: 1) Provide a comprehensive budgeting experience. 2) Exhibit the use of Excel in c
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!

Your question is very long and as per HOMEWORKLIB POLICY I am not allowed to solve questions with so many subparts. Still I have attached all the budgets till Income Statement. Request you to raise another question for Sensitivity Analysis.

Managerial Accounting
Sales Budget January February March April May Note
Projected Sales units           13,402.00           45,819.00           44,164.00         53,722.00         52,482.00 A
Sell price per unit                   44.00                   44.00                   44.00                 44.00                 44.00 B
Anticipated Sales Dollars         589,688.00      2,016,036.00      1,943,216.00 2,363,768.00 2,309,208.00 C=A*B
Total Selling & admin expenses Budget January February March
Sales in units            13,402.00            45,819.00            44,164.00 See A
Variable Selling & admin expenses per unit                   16.00                   16.00                   16.00 D
Anticipated Variable costs         214,432.00         733,104.00         706,624.00 E=A*D
Fixed Selling & admin cost              5,519.00              5,519.00              5,519.00 F
Total Anticipated Selling & admin cost         219,951.00         738,623.00         712,143.00 G=E+F
Budgeted Production in Units January February March April May
Sales in units            13,402.00            45,819.00            44,164.00         53,722.00         52,482.00 See A
Add: Desired Ending Inventory            27,033.21            26,056.76            31,695.98         30,964.38 H= 59% of A of next month. Because Desired Ending Inventory is 59% of next month sales units.
Total units needed           40,435.21           71,875.76           75,859.98         84,686.38
Less: Beginning Inventory                          -              27,033.21            26,056.76         31,695.98 I= 59% of A of same month. Because company is keeping 59% of current month sales as opening stock and January stock is 0 as mentioned in question.
Budgeted Production in Units           40,435.21           44,842.55           49,803.22         52,990.40 J
Material requirement Budget January February March April
Units to be produced           40,435.21           44,842.55           49,803.22         52,990.40 See J
Material required per unit                   11.00                   11.00                   11.00                 11.00 K
Total Production need         444,787.31         493,268.05         547,835.42       582,894.40 L=J*K
Add: Desired Ending Inventory         318,651.16         353,901.68         376,549.78                        -   M= 64.6% of L of next month. Because Desired Ending Inventory is 64.6% of next month production units.
Total Materials needed         763,438.47         847,169.73         924,385.20
Less: Beginning Inventory                          -           318,651.16         353,901.68       376,549.78 N= 64.6% of L of same month. Because company is keeping 64.6% of current month production as opening stock and January stock is 0 as mentioned in question.
Total materials needed to be purchased         763,438.47         528,518.57         570,483.52 O
Cost per unit of material                   23.00                   23.00                   23.00 P
Total materials cost 17,559,084.82 12,155,927.13 13,121,120.98 Q=P*O
Direct Labor Budget January February March
Number of Units to be produced           40,435.21           44,842.55           49,803.22 See J
Labor Hour required per unit                      1.00                      1.00                      1.00 R
Total Labor Hour needed           40,435.21           44,842.55           49,803.22 S=J*R
Labor Rate per Hour                   21.00                   21.00                   21.00 T
Total Direct Labor cost         849,139.41         941,693.55      1,045,867.62 U=S*T
Labor Hour required per unit                      1.00 See R
Labor Rate per Hour                   21.00 See T
Direct Labor cost per product                   21.00 V= R*T
Manufacturing overhead Budget January February March
Production in units           40,435.21           44,842.55           49,803.22 See J
Variable OH rate per unit                      5.00                      5.00                      5.00 W
Anticipated VOH costs         202,176.05         224,212.75         249,016.10 X= J*W
Fixed OH costs              5,625.00              5,625.00              5,625.00 Y
Total Anticipated OH costs         207,801.05         229,837.75         254,641.10 Z= X+Y
Predetermined OH Rate
Total estimated OH costs (for the quarter)         692,279.90                      5.12 Per DLH
Total estimated DLHs (for the quarter)         135,080.98
MOH per unit
Predetermined MOH Rate                      5.12
DLHs per unit                      1.00
MOH per unit                      5.12 AA
Product Cost Per unit
Direct Materials                   23.00 See P
Direct Labor                   21.00 See T
Manufacturing Overhead                      5.12 See AA
Total Unit Cost                   49.12 AB
Projected Income Statement January February March
Sales in Units           13,402.00           45,819.00           44,164.00 See A
Sales Revenue         589,688.00      2,016,036.00      1,943,216.00 See C
Cost of Good sold
Direct Materials         308,246.00      1,053,837.00      1,015,772.00 This is P*A
Direct Labor         281,442.00         962,199.00         927,444.00 This is T*A
Manufacturing Overhead            68,684.25         234,818.94         226,337.19 This is AA*A
Gross Profit (margin)         (68,684.25)       (234,818.94)       (226,337.19) AC
Selling & admin (operating) expenses         219,951.00         738,623.00         712,143.00 See G
Operating Income       (288,635.25)       (973,441.94)       (938,480.19) AD=AC-G